Mark had a board 25 1/2 inches long which he cut into 10 pieces. How long would the board be if it had 17 pieces of the same thickness?
a. 42 7/20 inches
b. 43 7/20 inches
c. 45 7/20 inches
d. 41 7/20 inches
I worked this one also and came up with Letter "A"
Please help...

Answer by Alan3354(69443)

Mark had a board 25 1/2 inches long which he cut into 10 pieces. How long would the board be if it had 17 pieces of the same thickness?
a. 42 7/20 inches
b. 43 7/20 inches
c. 45 7/20 inches
d. 41 7/20 inches
----------
25.5/10 = x/17
Cross multiply
10x = 25.5*17
x = 25.5*17/10
x = 43.35 inches
x = 43 7/20 inches